A 0.01kg bullet is fired horizontally into a 1 kg block of wood, initially at rest on a frictionless surface. After the impact, the block of wood with the bullet inside recoils with a velocity of 2m/s.

1. Is momentum conserved in this process??

2.Is energy conserved in this process?

3. What was the velocity of the bullet before impact with the block of wood?
Answer #1

Momentum is conserved in this process

As there is no heat losses incurred in the collision, Energy should be conserved

3) conservation of momentum

mu + 0 = (m + M)V

u = 1.01*2/0.01 = 202 m/s